OpenStudy (anonymous):

Pls just put the answer im really tired. Thank you very much!!! If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find ƒ(-1) + 5g(-3). -40 -38 7 2).If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find 2 · ƒ(4). 18 34 65 3).If ƒ(x ) = x 2 + 1 and g(x ) = 3x + 1, find [ƒ(4)] 2. 81 257 289

OpenStudy (jebonna):

dont be lazy xxxxxxx

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Pls help.

OpenStudy (academicgurusinc):

5g(-3) = 5[3(-3) + 1]= 5[-9 + 1]= 5(-8) = -40 Therefore, ƒ(-1) + 5g(-3) = -1 + (-40) = -41 Now try the other questions. We hope this helps. Regards, Academic GurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)

OpenStudy (anonymous):

 solve it @seancreed